OpenStudy (raden):

The cubic equation x^3 + Rx^2 + Sx + M = 0 has three different non zero roots R, S, and M. compute the value of the expression 1/S + 1/M

OpenStudy (3mar):

Hope that helps

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (raden):

i want get an exact value. @agent0smith any idea sir ?

OpenStudy (raden):

from the Vietas formula, get the equations : R + M + S = -R ... (1) RM + RS + MS = S ... (2) RMS = -S ... (3) or RM = -1

OpenStudy (raden):

so, the 2nd equation can be RS + MS = S + 1

OpenStudy (raden):

this is similar like x^3 + ax^2 + bx + c = 0 with a, b, and c are roots (non zero) a + b + c = -a or b + c = -2a ... (1) ab + ac + bc = b a(b + c) + bc = b ... (2) abc = -c ab = -1 a = -1/b ... (3) substitution (3) into (1), get b +c = -2(-1/b) = 2/b c = 2/b - b = (2 - b^2)/b and the 2nd equation : -1/b (b + 2/b - b) + b(2/b - b) = b -1/b (2/b) + 2 - b^2 = b -2/b^2 + 2 - b^2 = b times b^2 to both sides, get -2 + 2b^2 - b^4 = b^3 b^4 + b^3 - 2b^2 + 2 = 0

OpenStudy (raden):

by using horner diagram, get b^4 + b^3 - 2b^2 + 2 = 0 (b + 1)(b^3 - 2b + 2) = 0 b = -1 or the roots of b^3 - 2b + 2 = 0

OpenStudy (raden):

if b = -1 is satisfying this case. i get a = 1 and c = - 1. of couse it is not allowed , bcz the roots must be diferent

OpenStudy (raden):

and finally i think we must see the last solution from b^3 - 2b + 2 = 0

OpenStudy (raden):

is this allowed ? b + c = -2a = -2(-1/b) = 2/b ---> or c = (2/b - b) a(b + c) + bc = b bc = b - a(b + c) bc = b - (-1/b)(b + c) bc = b + (b+c)/b = b + (b + 2/b - b)/b = b + 2/b^2 = (b^3 + 2)/b^2 1/b + 1/c = (b + c)/(bc) = (2/b) / (b^3 + 2)/b^2 = 2b/ (b^3 + 2) = 2b /2b = 1

OpenStudy (raden):

but i cant give a proving that a , b, and c are not same :v